# Setting up for failure: automatic discovery of the neural mechanisms of cognitive errors

### v1

@@ 4 Results
Figure 4: Behaviourally realistic, but not task optimal, networks capture neural signatures of VWM. A $\bm{m}_{T}$ averaged over many accurate trials with distractor report feature varying, for trials with close (top) and far (bottom) stimulus distance, before (left, planes misaligned) and after (right, planes aligned) cue exposure. Colours represent report feature of cued (and accurately recalled) stimulus. Representations drawn from network in E (right), which best matches neural data (see below). B This qualitatively matches equivalent representations in macaque cortex - in this case there are only two possible cue values. We find a similar pre- and post-cue geometry for our index-cued networks, which are the closest analogue to the real task. C The real data (dots) further shows that this planar alignment (quantified as cosine similarity between normals) increases before and during cue exposure (grey stripe), and remains high throughout the second delay until time of recall. This is matched for our index cued networks when trained with a data-matched proportion of swap errors (5-10%). Even training with our DDPM-based method with a no-swap target distribution does not achieve this rapid increase during the cue period. A longer second delay period was chosen to encompass all experimental conditions in the original study. D Plane alignment over time for the feature-cued networks. Axes borders indicate the model used, as in Figure 3 . Importantly, only the model with distance-dependent swap probability matches the description of the biological data like its index-cued counterparts. This is quantified in Appendix C. Error bars show mean ± std across all different spatial configurations with the denoted stimulus location distance (see inset; purple is closest distractor, yellow is furthest). B C generated with data from Panichello2021
@@
We now compare the neural representations of stimuli during each delay to those of macacque lateral prefrontal cortex (LPFC) during a similar task [Panichello2021]. Here, two macacque monkeys performed a version of the multiitem delayed estimation task with fixed stimulus locations - one coloured square in the top and one in the bottom half of their field of vision. A binary visual cue at the point of fixation indicated which item is to be recalled. This task is most analogous to the index-cued version of our networks. As such, following subsections will begin by reproducing biological phenomena with these networks, before going on to make predictions about the role of the probe feature with our feature-cued networks. For the former, we match behavioural statistics; for the latter, we construct archetypal swap dependencies as the target distribution, to form predictions about neural representations.
@@
In the feature-cued case, where a continuous location is used to load and cue items from memory, this pattern is only captured when the network is trained with a location-dependent swapping rate (Figure 4D [right]), and not with a flat, probe feature-independent rate of swap errors. As motivated above, while we matched behavioural statistics for the index-cued task, which bears direct resemblance to the real task provided to monkeys, we modelled the form of the swap dependence off more sophisticated tasks [Emrich2012, Schneegans2017, McMaster2022] on arc in this task in order to make predictions about their neural representations. Furthermore, this network bears higher pre-cue representational alignment for distractors with closer probe feature values. This provides a prediction for the neural substrate of delay-time processes leading to memory retrieval error at cueing, previously described in the psychophysics literature as noising or drift in the abstract representations of stimuli along the probe dimension over time [Schneegans2017, McMaster2022]. Concretely, we predict that trial-by-trial decomposed analysis of the alignment between stimulus representations will reveal that swap errors occur more often when pre-cue representational alignment of colour is higher. This could be due to probe similarity, as in the case presented, or other factors such as attention before cueing in the fixed item location case.
@@
Item misselection   The previous subsection made the prediction that physically closer, hence more swap-prone, distractors cause higher report feature representation alignment before cue onset (Figure 4D [right]). We now consider the neural signatures of the retrieval errors at cueing time, which lead to swap errors. For the same macaque task, experimentalists later showed that item misselection is the primary cause of swap errors (Figure 5A; Alleman2024). The misselection is operationalised as follows: (i) before cue onset, items are stored as conjunctions between their probe (location - or just upper and lower in the experimental case) and report feature (colour) values (Figure 5A [left]); after cue onset, items are stored as conjunctions between their role (target vs. distractor) and report feature value (Figure 5A [right]) These are dubbed the spatial and report subspaces respectively; item ‘misselection’ is misprojection between these subspaces at cueing time. Swap errors can also arise at cueing time due to misinterpretation of the cue variable. Alternatively, report and probe features can be misbound prior to cueing. Alleman2024 find the least evidence for this mechanism, instead asserting that swap errors arise during or after cueing time. We now complement existing evidence of this neural basis of swap errors, which depends on across-trial statistical modelling of neural responses Alleman2024, with the representational geometry learned by our network.
@@
First, we apply the same linear decoding analysis as Alleman2024 to the memory subspace neural activity of our index cued network. These networks hold direct analogy to the task analysed by this work, so results can be directly compared. This method is based on training linear decoders on non-swap trials, then projecting neural activity onto the line connecting these nominal representations to representations that arise due to misbinding, cue misinterpretation, or item misselection. We defer details of this analysis to Alleman2024. As with the real macaque data, this analysis suggests that swap errors arise in our network during the second delay, rather than a misbound representations of memoranda Schneegans2017, McMaster2022.
@@
Next, we investigate our probe-cued networks for the existence of report subspace, where colours are bound to roles (target or distractor) with locations factored out, and the neural signature of swap errors during delay 2. Figure 5C shows a low-dimensional projection of $\bm{m}_{T}$ for a fixed set of locations and one fixed colour forming two parallel rings, coloured by the free stimulus. In the left column, when the fixed colour stimulus is cued, the bottom ring (negative PC3) is made up of the representation of swap trials, which are coloured by the uncued but mistakenly recalled colour, and the top ring (positive PC3) is coloured by the distractor colour when the fixed colour item is accurately recalled. Vice versa when the fixed colour stimulus is uncued (right column). The normal to these rings, which discriminates accurate and swap trials (Figure 5D [bottom]), is independent of the stimulus locations, only depending on the fixed stimulus colour considered (Figure 5D [top]). This makes this two ring structure a candidate for the geometry of the postcue report subspace [Alleman2024], which binds stimulus colour (around the ring) to its recalled role (choice of ring), and is independent of the original colour-location binding (Figure 5A). We have not considered exactly how this report subspace varies across different fixed report stimuli, just that the normals of these rings is shared whenever one of the items has a fixed colour, regardless of the item locations or whether the fixed item was cued (Figure 5C [left]). However, their alignment depends on colour similarity (Appendix E), suggesting an intuitive toroidal formation.
@@
Figure 5: RNNs trained with swap errors provide a prediction about the geometry of item misselection. A prior works [Alleman2024] suggest 3 causes of swap errors. Only misbinding errors arise during (or before) delay 1, prior to cueing. Misselection and misinterpretation cause incorrect transfer from a probe-report binding in delay 1 to a role-report binding in delay 2. B We repeat linear decoder analyses from Alleman2024 for our index-cued networks. Like the real data, these provide most evidence for swap errors arising during delay 2, ruling out misbinding errors. We defer full details to Alleman2024 C Trial-by-trial $\bm{m}_{T}$, projected to a shared set of principal components (PCs). One item’s colour is fixed to purple, and the other item’s colour is varied and used to colour the scatterplots. Left (right) axes show when the purple item is cued (a distractor), so colours indicate the uncued (target) colour. Probe features are fixed to a close (far) spatial configuration in the top (bottom) row, causing more (fewer) swap errors. Neural activities form two rings (with a gap in each due to the minimum margin between colours) - see main text. To illustrate this better, we used a network that swaps more often than seen for macaques (highest orange line in Figure 3A). D Left: there is high pairwise alignment in the third PC across different stimulus spatial configurations (i.e. fixed locations) if the fixed colour used to generate the two ring geometry is the same between these configurations - less so if the fixed colour is different (here, the opposite) for all pairs of spatial configurations. See Appendix E for a more graded pairwise comparison. This also applies to the first two PCs, unshown. Right: this third PC discriminates swapped versus correct trials in cases where the fixed colour item (purple here) is cued, across many different probe values.
@@ 5 Discussion
Our approach bridges the gap between flexibly modelling neural circuits and producing rich descriptions of behaviour, that were previously only achieved separately by task-optimised networks and abstract cognitive models, respectively. We did so by reversing the typical sequence of normative modelling - by directly training networks to reproduce the full distribution of behavioural responses, including suboptimalities, we bypass the need for an iterative, potentially heuristic set of ablations. These principles are most tightly shared by tiny RNNs [JiAn2025]. However these small networks, like previous approaches fitting to behavioural data [Xue2024] have only been used for discrete behaviour, e.g. categorical choices, for which simple maximum likelihood fitting to behavioural datasets suffices. Our method provides a novel, principled approach to holistically fitting to all, continuous dimensions of data. The neural predictions we made by reverse engineering behaviour in this way await experimental verification, but demonstrate the value of training networks to reproduce behavioural errors rather than optimising for performance.
@@
One limitation of our approach is its dependence on a generative behavioural model to produce training data, which may not be available for all tasks of interest. However, for many well-studied paradigms such models already exist, and recent advances in descriptive modelling [Binz2024] suggest that capturing the relevant statistical structure of behaviour, rather than normative or mechanistic correctness, is often sufficient and increasingly achievable.
@@
The present work can be extended in multiple directions. We used this VWM task as a testbed for our novel methodology because of its inherent multimodality, but our method can be extended to any other cognitive task associated with complex behavioural suboptimalities, or other non-trivial response distributions, such as reaction times [PardoVazquez2019]. For example, we report our success in training a model to perform a cue combination (multisensory integration; wei_cue_combination) task, where the network is trained to generate samples from the posterior of a Bayesian ideal observer. This task currently acts as a demonstration of the flexibility of our technique, rather than an extension to the mechanistic hypotheses it generates. Future work can extend in this direction.
@@
Within swap error analysis, our analysis of the representational geometry of stimuli and responses (Figures 4 and 5) mostly considers stationary snapshots of the population activity. Besides this, Figures 4D, E aggregate information across many trials. Future work should include trial-by-trial dynamical analysis. For example, the misselection process described in Figure 5A may be due to a failure in information loading [Stroud2023] during the cue period, which is made more likely due to the higher representational alignment we uncovered here. Better interpreting why RNNs capture neural phenomena only when trained explicitly to exhibit errors, beyond the descriptive summary we have provided here, can help refine this inquiry. Looking forward, there exists a much richer set of behavioural phenomena associated with VWM which can be modelled with our method. Further dependence of swap errors on set size [Bays2009, Emrich2012] and presentation order [vanEde2021, Gorgoraptis2011], and other, unimodal, response distributions which cannot be captured by moment-matching [Fritsche2020] are all suitable next paradigms to benefit from our method.
@@ Appendix A Bayesian Non-parametric model of Swap errors (BNS)
Stimulus array $Z=\{\bm{z}^{(n)}:n=1,2\}$ is denoted by the vector of pprobe (location) and rreport (colour) feature values: $\bm{z}^{(n)}=[z_{p}^{(n)},z_{r}^{(n)}]\in S^{1}\times S^{1}$. Item $n^{*}$ is cued after the first delay. The probe-dependent BNS Radmard2025 generative process defines a mixture distribution over estimates (recalled colour, $y$), based on the circular distance of the distractor from the cued item (i.e. in the probe dimension). The so-called swap function $f$ dictates the form of this dependence. For two items it is:
@@
To generate synthetic data, we fix one target swap function $f^{*}$, and perform the generative process up to component $\beta$, before embedding the samples into $\mathbb{R}^{2}$ and adding noise there, rather than using a circular emission distribution $p_{\epsilon}$. We also generate both feature values with a minimum margin of $\pi/4$, akin to previous multiitem tasks Schneegans2017, McMaster2022:
@@
Figure 6: Graphical model of BNS, adapted from Radmard2025. $\theta$ contains parameters for emission distribution $p_{\varepsilon}$. $\omega$ contains GP prior parameters and $\tilde{\pi}_{\text{correct}}$
@@
and use variational inference to fit $q(f)$, which approximates the dataset posterior over swap function $f$, which is equipped with a Gaussian process gps prior $p_{f}$. We defer details to Radmard2025. We do not consider the uniform component used in the introductory work, given the rarity of RNN behavioural samples away from the target modes.
@@ Appendix B Architecture and task details
Our RNN models a population of $n$ densely connected cortical pyramidal neurons Lyo2024. Each neuron integrates inputs from all other neurons in the population via a dendritic tree. Distal tree nodes are provided with sensory information, and all nodes are provided with a representation of time along the trial, necessary for training it as a diffusion model.
@@
where $k$ indexes over neurons, $\bm{i}$ indexes a particular path of nodes in the dendritic tree, and $j$ enumerates distal nodes connecting to $\bm{i}$. The population’s $n$ neurons [^2] Undeclared numerical constants are defined and provided values in Appendix G receive inputs from a dendritic tree with $L$ layers, each with a branching factor of $B_{l}$ from the previous layer. This results in each neuron axon synapsing onto $n\prod_{l=1}^{L}B_{l}$ dendritic leaf nodes. The rate vector $\bm{r}$ is projected via one set of weights $W_{r}\in\mathbb{R}^{nB_{1}\ldots B_{L}\times n}$, then used as inputs to the most distal layer, instead of $h$. As mentioned in the main text, this architecture choice was partly motivated by minimum requirements of expressivity. We started with standard architectures employing linear projections followed by point non-linearities (e.g., Yang2019), but we consistently struggled to achieve multimodal dynamics. A two-layer network at $\mathcal{F}$ could produce multimodality, but this is biologically implausible as it implies an intermediary neural population with instantaneous responses (zero time constant). We therefore adopted dendritic tree architectures [Lyo2024], which constrain a deep network’s weights to block-diagonal form, producing tree-like connectivity that reflects cortical pyramidal neuron morphology whilst enabling the required distributional flexibility. While a dendritic structure is not necessary for our method, it is a proven [Lyo2024], expressive structure that maintains plausibility in this regard.
Each node is also provided with a bias term. Before the cue offset ($t<0$), when the network is still accumulating information, this is a constant bias $b$. After cue offset, when the network is denoising during the WM second delay, this becomes a smoothly changing $n_{t}$-dimensional representation of time Ho2020, 10.5555/3295222.3295349 projected to a scalar with a unique set of weights for each dendritic node.
@@
To provide a sensory input to the ‘feature-cued’ network, we represent these stimuli with activations of a palimpsest conjunctively tuned neural population (Figure 7A [second row]; Matthey2015, Schneegans2017), then combine the tuning curve information before projection to the neural population using two sets of weights $W_{p},W_{r}$:

### v2

@@ 4 Results
Figure 4: Behaviourally realistic, but not task optimal, networks capture neural signatures of VWM. A $\bm{m}_{T}$ averaged over many accurate trials with distractor report feature varying, for trials with close (top) and far (bottom) stimulus distance, before (left, planes misaligned) and after (right, planes aligned) cue exposure. Colours represent report feature of cued (and accurately recalled) stimulus. Representations drawn from network in E (right), which best matches neural data (see below). B This qualitatively matches equivalent representations in macaque cortex - in this case there are only two possible cue values. We find a similar pre- and post-cue geometry for our index-cued networks, which are the closest analogue to the real task. C The real data (dots) further shows that this planar alignment (quantified as cosine similarity between normals) increases before and during cue exposure (grey stripe), and remains high throughout the second delay until time of recall. This is matched for our index cued networks when trained with a data-matched proportion of swap errors (5-10%). Even training with our DDPM-based method with a no-swap target distribution does not achieve this rapid increase during the cue period. A longer second delay period was chosen to encompass all experimental conditions in the original study. D Plane alignment over time for the feature-cued networks. Axes borders indicate the model used, as in Figure 3 . Importantly, only the model with distance-dependent swap probability matches the description of the biological data like its index-cued counterparts. This is quantified in Appendix C. Error bars show mean ± std across all different spatial configurations with the denoted stimulus location distance (see inset; purple is closest distractor, yellow is furthest). B C generated with data from Panichello and Buschman [2021]
@@
We now compare the neural representations of stimuli during each delay to those of macacque lateral prefrontal cortex (LPFC) during a similar task [Panichello and Buschman, 2021]. Here, two macacque monkeys performed a version of the multiitem delayed estimation task with fixed stimulus locations - one coloured square in the top and one in the bottom half of their field of vision. A binary visual cue at the point of fixation indicated which item is to be recalled. This task is most analogous to the index-cued version of our networks. As such, following subsections will begin by reproducing biological phenomena with these networks, before going on to make predictions about the role of the probe feature with our feature-cued networks. For the former, we match behavioural statistics; for the latter, we construct archetypal swap dependencies as the target distribution, to form predictions about neural representations.
@@
In the feature-cued case, where a continuous location is used to load and cue items from memory, this pattern is only captured when the network is trained with a location-dependent swapping rate (Figure 4D [right]), and not with a flat, probe feature-independent rate of swap errors. As motivated above, while we matched behavioural statistics for the index-cued task, which bears direct resemblance to the real task provided to monkeys, we modelled the form of the swap dependence off more sophisticated tasks [Emrich and Ferber, 2012, Schneegans and Bays, 2017, McMaster et al., 2022] on arc in this task in order to make predictions about their neural representations. Furthermore, this network bears higher pre-cue representational alignment for distractors with closer probe feature values. This provides a prediction for the neural substrate of delay-time processes leading to memory retrieval error at cueing, previously described in the psychophysics literature as noising or drift in the abstract representations of stimuli along the probe dimension over time [Schneegans and Bays, 2017, McMaster et al., 2022]. Concretely, we predict that trial-by-trial decomposed analysis of the alignment between stimulus representations will reveal that swap errors occur more often when pre-cue representational alignment of colour is higher. This could be due to probe similarity, as in the case presented, or other factors such as attention before cueing in the fixed item location case.
@@
Item misselection   The previous subsection made the prediction that physically closer, hence more swap-prone, distractors cause higher report feature representation alignment before cue onset (Figure 4D [right]). We now consider the neural signatures of the retrieval errors at cueing time, which lead to swap errors. For the same macaque task, experimentalists later showed that item misselection is the primary cause of swap errors (Figure 5A; Alleman et al. [2024]). The misselection is operationalised as follows: (i) before cue onset, items are stored as conjunctions between their probe (location - or just upper and lower in the experimental case) and report feature (colour) values (Figure 5A [left]); after cue onset, items are stored as conjunctions between their role (target vs. distractor) and report feature value (Figure 5A [right]) These are dubbed the spatial and report subspaces respectively; item ‘misselection’ is misprojection between these subspaces at cueing time. Swap errors can also arise at cueing time due to misinterpretation of the cue variable. Alternatively, report and probe features can be misbound prior to cueing. Alleman et al. [2024] find the least evidence for this mechanism, instead asserting that swap errors arise during or after cueing time. We now complement existing evidence of this neural basis of swap errors, which depends on across-trial statistical modelling of neural responses Alleman et al. [2024], with the representational geometry learned by our network.
@@
First, we apply the same linear decoding analysis as Alleman et al. [2024] to the memory subspace neural activity of our index cued network. These networks hold direct analogy to the task analysed by this work, so results can be directly compared. This method is based on training linear decoders on non-swap trials, then projecting neural activity onto the line connecting these nominal representations to representations that arise due to misbinding, cue misinterpretation, or item misselection. We defer details of this analysis to Alleman et al. [2024]. As with the real macaque data, this analysis suggests that swap errors arise in our network during the second delay, rather than a misbound representations of memoranda Schneegans and Bays [2017], McMaster et al. [2022].
@@
Next, we investigate our probe-cued networks for the existence of report subspace, where colours are bound to roles (target or distractor) with locations factored out, and the neural signature of swap errors during delay 2. Figure 5C shows a low-dimensional projection of $\bm{m}_{T}$ for a fixed set of locations and one fixed colour forming two parallel rings, coloured by the free stimulus. In the left column, when the fixed colour stimulus is cued, the bottom ring (negative PC3) is made up of the representation of swap trials, which are coloured by the uncued but mistakenly recalled colour, and the top ring (positive PC3) is coloured by the distractor colour when the fixed colour item is accurately recalled. Vice versa when the fixed colour stimulus is uncued (right column). The normal to these rings, which discriminates accurate and swap trials (Figure 5D [bottom]), is independent of the stimulus locations, only depending on the fixed stimulus colour considered (Figure 5D [top]). This makes this two ring structure a candidate for the geometry of the postcue report subspace [Alleman et al., 2024], which binds stimulus colour (around the ring) to its recalled role (choice of ring), and is independent of the original colour-location binding (Figure 5A). We have not considered exactly how this report subspace varies across different fixed report stimuli, just that the normals of these rings is shared whenever one of the items has a fixed colour, regardless of the item locations or whether the fixed item was cued (Figure 5C [left]). However, their alignment depends on colour similarity (Appendix E), suggesting an intuitive toroidal formation.
@@
Figure 5: RNNs trained with swap errors provide a prediction about the geometry of item misselection. A prior works [Alleman et al., 2024] suggest 3 causes of swap errors. Only misbinding errors arise during (or before) delay 1, prior to cueing. Misselection and misinterpretation cause incorrect transfer from a probe-report binding in delay 1 to a role-report binding in delay 2. B We repeat linear decoder analyses from Alleman et al. [2024] for our index-cued networks. Like the real data, these provide most evidence for swap errors arising during delay 2, ruling out misbinding errors. We defer full details to Alleman et al. [2024] C Trial-by-trial $\bm{m}_{T}$, projected to a shared set of principal components (PCs). One item’s colour is fixed to purple, and the other item’s colour is varied and used to colour the scatterplots. Left (right) axes show when the purple item is cued (a distractor), so colours indicate the uncued (target) colour. Probe features are fixed to a close (far) spatial configuration in the top (bottom) row, causing more (fewer) swap errors. Neural activities form two rings (with a gap in each due to the minimum margin between colours) - see main text. To illustrate this better, we used a network that swaps more often than seen for macaques (highest orange line in Figure 3A). D Left: there is high pairwise alignment in the third PC across different stimulus spatial configurations (i.e. fixed locations) if the fixed colour used to generate the two ring geometry is the same between these configurations - less so if the fixed colour is different (here, the opposite) for all pairs of spatial configurations. See Appendix E for a more graded pairwise comparison. This also applies to the first two PCs, unshown. Right: this third PC discriminates swapped versus correct trials in cases where the fixed colour item (purple here) is cued, across many different probe values.
@@ 5 Discussion
Our approach bridges the gap between flexibly modelling neural circuits and producing rich descriptions of behaviour, that were previously only achieved separately by task-optimised networks and abstract cognitive models, respectively. We did so by reversing the typical sequence of normative modelling - by directly training networks to reproduce the full distribution of behavioural responses, including suboptimalities, we bypass the need for an iterative, potentially heuristic set of ablations. These principles are most tightly shared by tiny RNNs [Ji-An et al., 2025]. However these small networks, like previous approaches fitting to behavioural data [Xue et al., 2024] have only been used for discrete behaviour, e.g. categorical choices, for which simple maximum likelihood fitting to behavioural datasets suffices. Our method provides a novel, principled approach to holistically fitting to all, continuous dimensions of data. The neural predictions we made by reverse engineering behaviour in this way await experimental verification, but demonstrate the value of training networks to reproduce behavioural errors rather than optimising for performance.
@@
One limitation of our approach is its dependence on a generative behavioural model to produce training data, which may not be available for all tasks of interest. However, for many well-studied paradigms such models already exist, and recent advances in descriptive modelling [Binz et al., 2025] suggest that capturing the relevant statistical structure of behaviour, rather than normative or mechanistic correctness, is often sufficient and increasingly achievable.
@@
The present work can be extended in multiple directions. We used this VWM task as a testbed for our novel methodology because of its inherent multimodality, but our method can be extended to any other cognitive task associated with complex behavioural suboptimalities, or other non-trivial response distributions, such as reaction times [Pardo-Vazquez et al., 2019]. For example, we report our success in training a model to perform a cue combination (multisensory integration; Beierholm et al. [2007]) task, where the network is trained to generate samples from the posterior of a Bayesian ideal observer. This task currently acts as a demonstration of the flexibility of our technique, rather than an extension to the mechanistic hypotheses it generates. Future work can extend in this direction.
@@
Within swap error analysis, our analysis of the representational geometry of stimuli and responses (Figures 4 and 5) mostly considers stationary snapshots of the population activity. Besides this, Figures 4D, E aggregate information across many trials. Future work should include trial-by-trial dynamical analysis. For example, the misselection process described in Figure 5A may be due to a failure in information loading [Stroud et al., 2023] during the cue period, which is made more likely due to the higher representational alignment we uncovered here. Better interpreting why RNNs capture neural phenomena only when trained explicitly to exhibit errors, beyond the descriptive summary we have provided here, can help refine this inquiry. Looking forward, there exists a much richer set of behavioural phenomena associated with VWM which can be modelled with our method. Further dependence of swap errors on set size [Bays et al., 2009, Emrich and Ferber, 2012] and presentation order [van Ede et al., 2021, Gorgoraptis et al., 2011], and other, unimodal, response distributions which cannot be captured by moment-matching [Fritsche et al., 2020] are all suitable next paradigms to benefit from our method.
@@ Appendix A Bayesian Non-parametric model of Swap errors (BNS)
Stimulus array $Z=\{\bm{z}^{(n)}:n=1,2\}$ is denoted by the vector of pprobe (location) and rreport (colour) feature values: $\bm{z}^{(n)}=[z_{p}^{(n)},z_{r}^{(n)}]\in S^{1}\times S^{1}$. Item $n^{*}$ is cued after the first delay. The probe-dependent BNS Radmard et al. [2025] generative process defines a mixture distribution over estimates (recalled colour, $y$), based on the circular distance of the distractor from the cued item (i.e. in the probe dimension). The so-called swap function $f$ dictates the form of this dependence. For two items it is:
@@
To generate synthetic data, we fix one target swap function $f^{*}$, and perform the generative process up to component $\beta$, before embedding the samples into $\mathbb{R}^{2}$ and adding noise there, rather than using a circular emission distribution $p_{\epsilon}$. We also generate both feature values with a minimum margin of $\pi/4$, akin to previous multiitem tasks Schneegans and Bays [2017], McMaster et al. [2022]:
@@
Figure 6: Graphical model of BNS, adapted from Radmard et al. [2025]. $\theta$ contains parameters for emission distribution $p_{\varepsilon}$. $\omega$ contains GP prior parameters and $\tilde{\pi}_{\text{correct}}$
@@
and use variational inference to fit $q(f)$, which approximates the dataset posterior over swap function $f$, which is equipped with a Gaussian process Rasmussen and Williams [2005] prior $p_{f}$. We defer details to Radmard et al. [2025]. We do not consider the uniform component used in the introductory work, given the rarity of RNN behavioural samples away from the target modes.
@@ Appendix B Architecture and task details
Our RNN models a population of $n$ densely connected cortical pyramidal neurons Lyo and Savin [2024]. Each neuron integrates inputs from all other neurons in the population via a dendritic tree. Distal tree nodes are provided with sensory information, and all nodes are provided with a representation of time along the trial, necessary for training it as a diffusion model.
@@
where $k$ indexes over neurons, $\bm{i}$ indexes a particular path of nodes in the dendritic tree, and $j$ enumerates distal nodes connecting to $\bm{i}$. The population’s $n$ neurons [^2] Undeclared numerical constants are defined and provided values in Appendix G receive inputs from a dendritic tree with $L$ layers, each with a branching factor of $B_{l}$ from the previous layer. This results in each neuron axon synapsing onto $n\prod_{l=1}^{L}B_{l}$ dendritic leaf nodes. The rate vector $\bm{r}$ is projected via one set of weights $W_{r}\in\mathbb{R}^{nB_{1}\ldots B_{L}\times n}$, then used as inputs to the most distal layer, instead of $h$. As mentioned in the main text, this architecture choice was partly motivated by minimum requirements of expressivity. We started with standard architectures employing linear projections followed by point non-linearities (e.g., Yang et al. [2019]), but we consistently struggled to achieve multimodal dynamics. A two-layer network at $\mathcal{F}$ could produce multimodality, but this is biologically implausible as it implies an intermediary neural population with instantaneous responses (zero time constant). We therefore adopted dendritic tree architectures [Lyo and Savin, 2024], which constrain a deep network’s weights to block-diagonal form, producing tree-like connectivity that reflects cortical pyramidal neuron morphology whilst enabling the required distributional flexibility. While a dendritic structure is not necessary for our method, it is a proven [Lyo and Savin, 2024], expressive structure that maintains plausibility in this regard.
Each node is also provided with a bias term. Before the cue offset ($t<0$), when the network is still accumulating information, this is a constant bias $b$. After cue offset, when the network is denoising during the WM second delay, this becomes a smoothly changing $n_{t}$-dimensional representation of time Ho et al. [2020], Vaswani et al. [2017] projected to a scalar with a unique set of weights for each dendritic node.
@@
To provide a sensory input to the ‘feature-cued’ network, we represent these stimuli with activations of a palimpsest conjunctively tuned neural population (Figure 7A [second row]; Matthey et al. [2015], Schneegans and Bays [2017]), then combine the tuning curve information before projection to the neural population using two sets of weights $W_{p},W_{r}$: